Referee of the article

“DIFFERENT GOALS IN MULTISCALE SIMULATIONS AND

HOW TO REACH THEM ” by Pierrick Tranouez and Antoine Dutot

The paper deals with different simulations using multiscale techniques. Three examples are chosen : maintaining a summary, using an environmental marker introducing an history in the data and using a knowledge on the behavior of the different scales in order to handle them at the same time. The paper is interesting and the discussion well carried out, but the text seems to have been hastily written. First, the loose style upsets the average reader : please avoid contractions which pertain to the spoken style, (1) and also exclamations like ”hey”. Second, some sentences are difficult to understand others are completely unintelligible (2). I suggest to accept the paper after a thorough revision.

(1) [AKNOWLEDGEMENTS] → [ACKNOWLEDGEMENTS] Now they´ re grown-ups ... . But they won´ t. Hey, we must start at some time. Comment : do not withdraw the aknowlegements however. Just write them in a more noble style

(2) In Conclusion : Our research group works on complex systems and focused on the computer representation of their hierarchical/holarchical characteristics ... & We know believe we start to have sound multiscale methods, and must focus on the realism of the applications,...